# DutyTTE: Deciphering Uncertainty in Origin-Destination Travel Time Estimation

### v1

@@ Introduction
First, baselines for travel time uncertainty quantification include: (1) T-WDR-MIS: Utilizes a transformer for path prediction as implemented in (Shi et al. 2024), with WDR (Wang, Fu, and Ye 2018b) for TTE. It uses prediction heads to output confidence intervals, trained with MIS loss (Gneiting and Raftery 2007). (2) MWSL-TTE (Wang et al. 2023): Sums estimated travel time across all segments of the predicted path. Uncertainty is based on variance from a learned Gaussian distribution for each segment. (3) DeepOD-MIS: Uses DeepOD (Yuan et al. 2020b) with prediction heads for confidence intervals, trained with MIS loss. (4) DOT-dropout: Incorporates MC Dropout (Zhu and Laptev 2017) into DOT (Lin et al. 2023), sampling 10 trajectories during inference. (5) DOT-MIS: Equips DOT (Lin et al. 2023) with prediction heads for confidence intervals, trained with MIS loss.
The desired confidence level of the confidence interval for travel time is set at $90\%$ (Angelopoulos et al. 2022; Bates et al. 2021). To evaluate uncertainty quantification, we use Interval Width (IW) and Prediction Interval Coverage Probability (PICP) (Lawless and Fredette 2005; Zhou et al. 2021), where PICP is defined as the proportion of ground truth values covered by the estimated intervals. Details about experimental settings are in the appendix.
Additionally, the following baselines are used for travel time estimation: (1) TEMP (Wang et al. 2019): Averages historical travel times with similar ODT. (2) GBDT (Chen and Guestrin 2016):Uses XGBoost for machine learning-based estimation. (3) ST-NN(Jindal et al. 2017): Jointly predicts travel distance and time.
We also compare methods for path prediction: (1) Dijkstra’s algorithm (DA)(Johnson 1973) Searches for the shortest path for given OD pairs. (2) NMLR (Jain et al. 2021): Learns segment weights and searches for paths with the largest weight at each step. (3) Key Segment (KS) (Tian et al. 2023): Detects a relay vertex for given OD pairs, predicting paths from origin to relay, then to the destination. (4) Transformer: Utilizes a transformer (Vaswani et al. 2017) for path prediction as implemented in (Shi et al. 2024). (5) GDP (Shi et al. 2024): uses a diffusion-based model to learn path patterns and predict paths via transformers.
Table 3: Performance of travel time estimation.
Datasets City A / City B
@@
TEMP 517.33/508.06 379.17/340.52 34.91/32.50
GBDT 309.45/372.10 234.03/271.81 33.98/31.90
STNN 311.80/317.48 246.67/244.26 33.28/32.93
DeepOD 336.51/329.74 263.73/245.12 31.07/30.12
MWSL-TTE 309.18/304.89 239.62/227.76 29.86/28.02
@@
Table 4: Performance of path prediction.
Datasets City A City B
@@
Improvement 7.48% 10.53% 5.85% 12.50%
@@
As shown in Table 1, DutyTTE excels in travel time uncertainty quantification.
MWSL-TTE and DOT-dropout underperform in PICP compared to methods like T-WDR-MIS, DeepOD-MIS, and DOT-MIS, which predict confidence intervals more effectively, highlighting the insufficiency of learned variances in covering ground truth during testing.
T-WDR-MIS surpasses DeepOD-MIS in PICP by leveraging recent travel time information in the predicted path, underscoring the importance of first predicting paths and then quantifying uncertainty based on travel time information of those paths.
While DOT-MIS and T-WDR-MIS lead to good results, they are limited by inaccuracies in path prediction. They fail to model travel time uncertainty of road segments under complex contexts, leading to inaccurate quantification of travel time uncertainty.
In contrast, DutyTTE predicts accurate paths and avoids using untraveled segment data by enhancing alignment with the ground truth by DRL techniques. It captures segment uncertainty under complex contexts using a MoEUQ, which adaptively selects experts for handling segment uncertainty in varied conditions. As a result, DutyTTE effectively learns the impact of travel time in each segment on overall uncertainty, achieving superior performance.
@@
MWSL-TTE and T-WDR are more accurate than DOT and DeepOD by predicting potential paths and utilizing recent travel time information. T-WDR outperforms MWSL-TTE; while MWSL-TTE uses weak supervision for segment travel times, T-WDR leverages recent data in road segments and its WDR architecture to capture correlations among influencing factors.
@@
Dataset City A
Metric RMSE ($\downarrow$) MAE (s) ($\downarrow$) MAPE (%) ($\downarrow$) PICP (%) ($\uparrow$) IW (s)
w/o-M 282.41 198.47 24.32 88.63 804.52
w/o-P 302.93 219.32 27.28 87.15 840.31
w/o-C 278.22 195.70 23.55 87.01 744.12
DutyTTE 278.22 195.70 23.55 92.12 773.88
Figure 3: Reward curve in the training process.
@@
To verify the effectiveness of components, we conduct an ablation study with the following DutyTTE variants: 1) w/o-M: replacing the mixture of experts guided uncertainty quantification mechanism with a LSTM. 2) w/o-P: removing the policy loss and using only cross-entropy loss in path prediction. 3) w/o-C: eliminating the calibration of predicted confidence intervals.
We compare the performance of these variants with DutyTTE, and the experimental results are presented in Table 5. We observe the following: 1) The mixture of experts guided uncertainty quantification mechanism is crucial for accurate uncertainty quantification, as MoEUQ effectively learns the uncertainty of each segment under different contexts. 2) Integrating metrics that measure the overall similarity between predicted paths and the ground truth into the policy loss during training leads to improvements in path prediction performance. 3) After calibration, the width of the estimated interval increases by 4.00%, while PICP increases by 5.87%. Additionally, the calibrated results achieve the desired PICP of 90%, demonstrating the effectiveness of the calibration method in providing statistical guarantees for the estimated confidence intervals. Details about the analysis of hyper-parameters are in the appendix.
@@
We evaluate the average execution time per batch, consisting of 128 samples from City A and City B, for efficiency validation, as shown in Figure 6. DutyTTE uses a DRL method for training; during prediction, the policy network allows for fast inferences. Given DutyTTE’s good performance, we see this as a promising trade-off between efficiency and effectiveness.
Table 6: Execution time comparison (seconds per batch).
Dataset DA DOT T-WDR-MIS MWSL-TTE GDP DutyTTE
City A 0.0722 1.362 0.2675 0.1289 6.916 0.2706
City B 0.0718 1.293 0.2649 0.1357 6.859 0.2682
@@ Conclusion
We introduce DutyTTE for quantifying uncertainty in OD travel times. Using a DRL approach, we optimize path prediction to align closely with the ground truth and enhance uncertainty quantification. To capture travel time uncertainty across road segments, we propose a mixture of experts mechanism that models the impact of segments in complex contexts. We also calibrate confidence intervals to ensure statistical guarantees. Experiments on two datasets validate DutyTTE’s effectiveness in path prediction, travel time uncertainty quantification.

### v2

@@ Introduction
First, baselines for travel time uncertainty quantification include: (1) T-WDR-MIS: Using a transformer for path prediction, with WDR (Wang, Fu, and Ye 2018b) for TTE. It uses prediction heads to output confidence intervals, trained with MIS loss. (2) MWSL-TTE (Wang et al. 2023): Sums estimated travel time across all segments of the predicted path. Uncertainty is based on variance from a learned Gaussian distribution for each segment. (3) DeepOD-MIS: Uses DeepOD (Yuan et al. 2020b) with prediction heads for confidence intervals, trained with MIS loss. (4) DOT-dropout: Incorporates MC Dropout (Zhu and Laptev 2017) into DOT (Lin et al. 2023). (5) DOT-MIS: Equips DOT with prediction heads for confidence intervals, trained with MIS.
The desired confidence level of the confidence interval for travel time is set at $90\%$ (Angelopoulos et al. 2022; Bates et al. 2021). To evaluate uncertainty quantification, we use Interval Width (IW) and Prediction Interval Coverage Probability (PICP) (Lawless and Fredette 2005; Zhou et al. 2021), where PICP is defined as the proportion of ground truth values covered by the estimated intervals.
Additional baselines for TTE: (1) TEMP (Wang et al. 2019): Estimate travel time using neighboring trips of a query. (2) GBDT (Chen and Guestrin 2016): A machine learning method. (3) ST-NN(Jindal et al. 2017): Jointly predicts travel distance and time.
We also compare methods for path prediction: (1) Dijkstra’s algorithm (DA)(Johnson 1973) Searches for the shortest path for given OD pairs. (2) NMLR (Jain et al. 2021): Learns segment weights and searches for paths with the largest weight at each step. (3) Key Segment (KS) (Tian et al. 2023): Detects a relay vertex for given OD pairs, predicting paths from origin to relay, then to the destination. (4) Transformer: Utilizes a transformer (Vaswani et al. 2017) for path prediction as implemented in (Shi et al. 2024). (5) GDP (Shi et al. 2024): uses a diffusion-based model to learn path distributions and predict paths via transformers.
Datasets Chengdu / Xi’an
@@
TEMP 317.33/312.06 259.17/249.62 32.16/32.07
GBDT 315.91/318.10 247.03/245.83 31.03/30.12
STNN 324.62/336.56 268.67/264.95 33.28/32.93
DeepOD 320.51/319.74 253.73/251.06 31.87/30.62
MWSL-TTE 307.18/304.89 239.62/227.76 29.86/27.58
@@
Table 3: Performance of travel time estimation.
Datasets Chengdu Xi’an
@@
Improvement 7.48% 10.53% 5.24% 12.50%
Table 4: Performance of path prediction.
@@
T-WDR-MIS and MWSL-TTE outperform DeepOD-MIS in PICP by leveraging estimated travel time in the predicted path, highlighting the importance of predicting paths before quantifying uncertainty. T-WDR-MIS predicts paths autoregressively by maximizing the likelihood of each segment independently, while MWSL-TTE estimates transition weights between segments to predict paths. Both methods suffer from error accumulation in path prediction, limiting their ability to provide reliable path information for uncertainty quantification.
In contrast, DutyTTE enhances alignment with ground truth using DRL techniques to predict accurate paths while avoiding untraveled segment data. It employs MoEUQ to capture segment-level uncertainty in complex contexts by adaptively selecting experts to handle varying conditions. This enables DutyTTE to effectively learn the contribution of each segment’s travel time to overall uncertainty, achieving superior performance.
@@
MWSL-TTE and T-WDR are more accurate than DOT and DeepOD by predicting potential paths and utilizing recent travel time information. T-WDR outperforms MWSL-TTE; while MWSL-TTE uses weak supervision for segment travel times, T-WDR leverages recent data in road segments and its wide-deep-recurrent architecture to capture correlations among influencing factors.
@@
Dataset Chengdu
Metric RMSE ($\downarrow$) MAE($\downarrow$) MAPE ($\downarrow$) PICP ($\uparrow$) IW
w/o-M 282.41 198.47 24.32 88.63 781.52
w/o-P 302.93 219.32 27.28 87.15 836.09
DutyTTE 278.22 195.70 23.55 91.02 763.65
@@
Figure 3: Reward curves during training.
@@
To verify the effectiveness of components, we conduct an ablation study with the following DutyTTE variants: 1) w/o-M: removing the mixture of experts guided uncertainty quantification mechanism. 2) w/o-P: removing the policy loss and using only cross-entropy loss in path prediction.
We compare the performance of these variants with DutyTTE, and the experimental results are presented in Table 5. We observe the following: 1) The MoEUQ is crucial for accurate uncertainty quantification, as MoEUQ effectively learns the uncertainty of each segment under different contexts. 2) Integrating metrics that measure the overall similarity between predicted paths and the ground truth into the policy loss during training leads to improvements in path prediction performance.
Hyper-parameter Analysis
We demonstrate the effectiveness of key hyper-parameters on testing sets. In Figure 4 (a), we set the value of $k$ and the number of experts as follows: $c_{1}:k=1,C=8$, $c_{2}:k=2,C=8$, $c_{3}:k=4,C=8$, $c_{4}:k=6,C=8$. In Figure 4 (b), we configure the policy loss function with the following hyper-parameters: $c_{5}:\omega=1,\beta=50,\gamma=1$, $c_{6}:\omega=1,\beta=100,\gamma=1$, $c_{7}:\omega=1,\beta=50,\gamma=0.5$, $c_{8}:\omega=0,\beta=0,\gamma=0$. We observe the following results.
1) In Figure 4 (a), setting the value of $k$ to 4 and the number of experts to $8$ can lead to optimal performance. This can be attributed to the fact that having too few $k$ and experts limits MoEUQ’s capability, as it can only handle a limited number of contexts for evaluating travel time uncertainty in a segment. However, if the number of experts is too large, the model complexity also increases, complicating the training process and compromising performance.
2) In Figure 4 (b), applying the policy loss function with different parameter settings, such as $c_{5}$, $c_{6}$, and $c_{7}$, results in improved performance compared to not using the policy loss function, as seen with $c_{8}$. Among these configurations, $c_{5}$ stands out as a better option for more accurate path prediction compared to $c_{6}$ and $c_{7}$. Achieving more accurate path predictions also leads to improvements in travel-time estimation and uncertainty quantification.
Figure 4: Effectiveness of hyper-parameters.
Analysis of the MoE guided uncertainty quantification
The experimental results in Figures 5 and Figure 6 demonstrate the effectiveness of the MoE method in quantifying uncertainty for road segment travel times. In Figure 5, 2000 road segment samples under various traffic conditions were randomly selected, with travel time variance calculated from historical trajectories observed 10 minutes before trip departure. This variance reflects the uncertainty of travel time for each segment. The t-SNE visualization reduces high-dimensional embeddings of these segments into two dimensions for comparison. The results demonstrate that with MoE, clusters of road segments with low and high travel time variance (indicating lower and higher uncertainty) are more distinct compared to the scenario without MoE, where cluster boundaries are less defined. This underscores MoE’s ability to better capture and differentiate travel time uncertainty under varying traffic conditions, improving overall uncertainty quantification. Additionally, Figure 6 shows the most frequently activated combination of experts for road segments with varying levels of uncertainty. The distribution of activated experts differs across uncertainty levels, indicating that the model dynamically employs diverse combinations of experts to effectively capture the uncertainty characteristics of road segments under different traffic conditions.
Dataset DA DOT T-WDR MWSL GDP DutyTTE
Chengdu 0.0722 1.362 0.2675 0.1289 5.093 0.2706
Xi’an 0.0718 1.293 0.2649 0.1357 4.896 0.2682
Table 6: Execution time comparison (seconds per batch).
Figure 5: t-SNE Visualization for embeddings of road segments under varying contexts.
Figure 6: Expert activation distributions.
@@
We evaluate the average execution time per batch, consisting of 128 samples from Chengdu and Xi’an, for efficiency validation, as shown in Table 6. DutyTTE uses a DRL method for training; during prediction, the policy network allows for fast inferences. Given DutyTTE’s good performance, we see this as a promising trade-off between efficiency and effectiveness.
@@ Conclusion
We introduce DutyTTE for quantifying uncertainty in OD travel times. Using a DRL approach, we optimize path prediction to align closely with the ground truth and enhance uncertainty quantification. To capture travel time uncertainty across road segments, we propose a mixture of experts mechanism that models the impact of segments in complex contexts. Experiments on two datasets validate DutyTTE’s effectiveness in path prediction, travel time uncertainty quantification.
Acknowledgment
This work was supported by the Fundamental Research Funds for the Central Universities (Grant No. 2023YJS030).